# Weight-Entanglement Meets Gradient-Based Neural Architecture Search

### v1

@@ 4.2 Cell-based search spaces
We now start our comparative analysis of single and two-stage approaches by applying them to cell-based spaces that serve as focal points in the single-stage NAS literature to evaluate TangleNAS against DrNAS and SPOS. Here, we use the popular NAS-Bench-201 (NB201) (Dong and Yang, 2020) and DARTS (Liu et al., 2019) search spaces. We refer the reader to Appendix C for details about these spaces and Section E.4 for the experimental setup.
Our contribution on these spaces is two-fold. First, we study the effects of weight-entanglement on cell-based spaces in conjunction with single-stage methods. To this end, we entangle the weights of similar operations with different kernel sizes on both search spaces. For NB201, the weights of the 1x1 and 3x3 convolutions are entangled, and in the DARTS search space, the weights of dilated and separable convolutions with kernel sizes 3x3 and 5x5 are entangled. Second, we study SPOS on the NB201 and DARTS search spaces. To the best of our knowledge, we are the first to study a two-stage method like SPOS in such cell search spaces.
@@
Search Type Optimizer Supernet CIFAR10 CIFAR100 ImageNet16-120 Search Time (hrs)
Single-Stage DrNAS WS $\mathbf{94.36\pm 0.00}$ $72.245\pm 0.732$ $\mathbf{46.37\pm 0.00}$ 20.9166
TangleNAS WE $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$ 20.3611
Two-Stage SPOS+RS WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.1460$ 29.51388
SPOS+ES WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.0149$ 26.74721
Table 3: Comparison of test-accuracies of single and two-stage methods on NB201 search space
Search Type Optimizer Supernet CIFAR10 ImageNet Search Time(hrs)
Single-Stage DrNAS WS $2.625\pm 0.075$ 26.29 9.12
TangleNAS WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.691}$ 7.44
Two-Stage SPOS+RS WE $2.965\pm 0.072$ 27.114 18.7444
SPOS+ES WE $3.200\pm 0.065$ 27.320 14.794
Table 4: Comparison of test errors of single and two-stage methods on the DARTS search space
@@ 4.3 Macro Search Spaces
Given the promising results of TangleNAS on the toy and cell-based spaces, we now extend our evaluation to the home base of two-stage methods. We study TangleNAS on a vision transformer space (AutoFormer-T) and a convolutional space (MobileNetV3), which have been proposed and explored using two-stage methods by Chen et al. (2021a) and Cai et al. (2020), respectively, as well as a language model transformer search space built around GPT-2 Radford et al. (2019).
@@ 4.3.1 AutoFormer
Search Type Optimizer CIFAR10 CIFAR100
Inherit Fine-tune Retrain Params FLOPS Inherit Fine-tune Retrain Params FLOPS
Single-Stage TangleNAS 93.57 $\mathbf{97.702\pm 0.017}$ $\mathbf{97.872\pm 0.054}$ 8.6847M 1.946 76.59 $\mathbf{82.615\pm 0.064}$ $\mathbf{82.668\pm 0.161}$ 8.64874M 1.9388G
Two-Stage SPOS+RS 94.29 $97.605\pm 0.038$ $97.767\pm 0.024$ 8.51161M 1.91G 78.210 $82.407\pm 0.026$ $82.210\pm 0.14242$ 8.47558M 1.9046G
SPOS+ES 94.10 $97.632\pm 0.047$ $97.6425\pm 0.023$ 7.230286M 1.6595 77.97 $82.517\pm 0.140$ $82.5175\pm 0.114$ 8.2447M 1.859G
Table 5: Evaluation on the AutoFormer-T space for CIFAR10 and CIFAR100.
We evaluate TangleNAS on the AutoFormer-T space introduced by Chen et al. (2021a), which is based on vision transformers. The search space consists of the choices of embedding dimensions and number of layers, and for each layer, its MLP expansion ratio and the number of heads it uses to compute attention. More details can be found in Table 19 in the appendix. The embedding dimension choice is held constant throughout the network while the number of heads and the MLP expansion ratio varies for every layer. This amounts to a total search space size of about $10^{13}$ architectures. We train our supernet using the same training hyperparameters as used in AutoFormer.
NAS Method SuperNet-Type ImageNet Datasets Params FLOPS
CIFAR10 CIFAR100 Flowers Pets Cars
SPOS+ES AutoFormer-T 75.474 98.019 86.369 98.066 91.558 91.935 5.893M 1.396G
TangleNAS AutoFormer-T 78.842 98.249 88.290 98.066 92.347 92.396 8.98108M 2.00G
SPOS+ES AutoFormer-S 81.700 99.10 90.459 97.90 94.8529 92.5447 22.9M 5.1G
TangleNAS AutoFormer-S 81.964 99.12 90.459 98.3257 95.07 92.3707 28.806M 6.019G
Table 6: Evaluation on the AutoFormer-T space on downstream tasks. ImageNet accuracies are obtained through inheritance, whereas accuracies for the other datasets are achieved through fine-tuning the imagenet-pretrained model.
Table 5 and Table 6 evaluates TangleNAS against AutoFormer on the AutoFormer-T space. Interestingly, we observe that although AutoFormer sometimes outperforms TangleNAS upon inheritance from the supernet, the TangleNAS architectures are always better upon fine-tuning and much better upon retraining. For ImageNet-1k we obtain an accuracy of $78.842$% by inheriting the weights of the optimal architecture in contrast to $75.474$% by unconstrained evolutionary search on the on AutoFormer-T space (Chen et al., 2021a), i.e. a net improvement of $3.368\%$ (see Table 6).
@@ 4.3.2 MobileNetV3
Search Type Optimizer Top-1 acc Params FLOPS
Single-Stage TangleNAS 77.424 7.58M 528.80M
Two-Stage OFA+RS 77.046 6.87M 369.16M
OFA+ES 77.050 7.21M 420.50M
Largest Arch - 77.336 7.66M 566.17M
Table 7: Evaluation on MobileNetV3
Next, we study a convolutional search space based on the MobileNetV3 architecture. The search space is defined in Table 19 in the appendix and contains about $2\times 10^{19}$ architectures. This follows from the search space designed by OFA (Cai et al., 2020), which searches for kernel-size, number of blocks and channel-expansion factor. Unlike OFA, we do not use the progressive shrinking scheme for the supernet training but activate all choices in our supernet at all times.
Table 7 shows that on this OFA search space, TangleNAS still outperforms OFA (based on both unconstrained evolutionary and random search on the pre-trained OFA supernet). Notably, TangleNAS even yields an architecture with higher accuracy than the largest architecture in the space (while OFA yields worse architectures).
4.3.3 Language Modelling Space
Finally, given the growing interest in efficiently training large language models and recent developments in scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), we study our efficient and scalable TangleNAS method on a language-model space.
We create our language model space around a smaller version of Andrej Karpathy’s nanoGPT model [^1] https://github.com/karpathy/nanoGPT. In this transformer search space, we search for the embedding dimension, number of heads, number of layers and MLP ratios as defined in Table 19 (in the appendix). For each of these transformer search dimensions we consider 3 choices. We evaluate our small language model on the TinyStories (Eldan and Li, 2023) dataset. We improve over the SPOS+ES and SPOS+RS baselines as presented in Table 8. This improvement is statistically very significant with a two-tailed p-value of $0.0064$ for ours v/s SPOS+ES and p-value of $0.0127$ for ours v/s SPOS+RS.
[…]
Single-Stage TangleNAS $\mathbf{1.412\pm 0.011}$ 4.104 87.01M 93.88s
Two-Stage SPOS+RS $1.433\pm 0.005$ 4.191 77.57M 85.169s
SPOS+ES $1.444\pm 0.013$ 4.238 78.75M 87.10s
Table 8: Comparison of single and two stage methods on language model search space. We report the test loss and perplexity on the tinystories dataset
@@ 5 Results and Discussion
For a more in-depth evaluation, we now compare different properties of single-stage and two-stage methods, starting from their any-time performance, impact of the portion of train-val split on their performance and centered-kernel alignment of the supernet feature maps. Moreover, we also study the effect of pretraining, fine-tuning and retraining on the AutoFormer space for CIFAR10 and CIFAR100 and downstream performance of imagenet pre-trained best model on several classification datasets. Finally, we discuss the insights derived from our single-stage approach in designing architectures on real world tasks.
@@
In practice, we observe that vanilla gradient-based NAS methods are memory and compute expensive in comparison to both two-stage methods and our TangleNAS approach with weight-superposition (described in Figure 2). In particular on the NB201 and DARTS search spaces we observe a 25.276% and 35.545% reduction in memory requirements for TangleNAS over DrNAS with WS. This issue only exacerbates for weight-entangled spaces like AutoFormer and MobileNetV3, making the application of vanilla gradient-based methods practically infeasible. This is because vanilla gradient-based methods incur an $\mathcal{O}(n)$ overhead over TangleNAS in terms of space and time complexity, where n is the number of operation choices.
Any-Time performance
An essential characteristic of any neural architecture search (NAS) method is the attainment of strong any-time performance. End users of NAS techniques frequently prioritize the expedited identification of competitive architectures within limited time constraints. This becomes particularly crucial in light of the escalating computational expenses associated with training sophisticated and large neural networks, such as Transformers. Strong any-time performance is a pivotal criterion for the effective applicability of NAS in such resource-intensive domains. We thus compare the any-time performance of TangleNAS. Figure 3 demonstrates that TangleNAS (DrNAS+WE) is faster than its baseline method DrNAS+WS. Similarly, Figure 4 shows that TangleNAS has better anytime performance than AutoFormer.
(a) CIFAR100
@@
Figure 3: Test accuracy evolution over epochs for NB201
(a) CIFAR10
(b) CIFAR100
Figure 4: Any Time performance curves of AutoFormer vs. Ours
@@
Effect of fraction of training data
One-shot NAS commonly employs a 50%-50% train-valid split for cell-based spaces and 80%-20% for weight-entangled spaces. To avoid potential confounding factors, we also study our method across different data splits for each search space. The results reported in Appendix A are largely similar across splits. In particular, we observe that single-stage methods are quite robust and performant across training fractions and search spaces, in comparison to two-stage methods.
@@
In transformer spaces reducing the mlp-ratio in the initial layers has a relatively low impact on performance (Figure 5) and can often work competitively or outperform handcrafted architectures. This observation is consistent across ViT and Language Model spaces. Conversely, number of heads and embedding dimension have a significant impact on the quantitative metric of choice. Pruning a few of the final layers also has a relatively low impact on performance. In the MobileNetV3 space, we find a strong preference for larger number of channels and larger network depth. In contrast, we discover that scaling laws for transformers may not necessary apply in convolutional spaces especially for kernel sizes. We observe that earlier layers favor 5x5 kernel sizes while later ones prefer 7x7 kernel sizes (3 being the smallest and 7 the largest).
@@
We study the impact of inheriting, fine-tuning, and retraining on the CIFAR10 and CIFAR100 datasets on the AutoFormer space. We observe that retraining almost always outperforms fine-tuning and inheriting. This brings to question the rank correlation between the inherited accuracy and retraining accuracy of two-stage methods and the training interference (Xu et al., 2022) in two-stage methods. Strong rank-correlation is desirable especially in two-stage methods as these methods rely on the performance proxy upon inheritance for the black-box search. Furthermore, we observe that while the SPOS+RS and SPOS+ES methods are performant upon inheritance, TangleNAS outperforms these models after fine-tuning and training from scratch. Improved supernet training properties for single-stage methods is further supported by Centered Kernel Alignment (CKA) analysis (Kornblith et al., 2019) (Table 25).
@@
Finally, we study the impact on fine-tuning the best model obtained from the search on downstream datasets. We follow the fine-tuning pipeline proposed in AutoFormer and fine-tune on different fine and coarse-grained datasets. We observe from Table 6 that the architecture discovered by TangleNAS on ImageNet is much more performant in fine-tuning to various datasets (CIFAR-10, CIFAR-100, Flowers, Pets and Cars) than the architecture discovered by SPOS.
6 Conclusion
In this paper, we investigate two of the most widely used NAS methods: single-stage and two-stage, each traditionally applied to different types of search spaces. We extend the applicability of single-stage NAS methods to weight-entangled spaces, a domain primarily explored by two-stage methods. By doing so, we aim to bridge the gap between these two distinct sub-fields. We empirically evaluate our proposed single-stage method, TangleNAS, on a diverse set of weight-entangled search spaces and tasks, showcasing its ability to outperform conventional two-stage NAS methods while enhancing search efficiency. Furthermore, we highlight the advantages of introducing weight-entanglement into traditional cell-based spaces, whenever feasible. Doing so not only improves the speed of single-stage methods but also reduces memory consumption. The modular nature of TangleNAS enables it to accommodate any differentiable optimizer and hence take advantage of the continually-improving set of methods for gradient-based NAS. Our results on various macro-level search spaces are already very encouraging and we therefore hope that this line of work will enable the NAS community to further improve modern architectures like Transformers.
6.0.1 Acknowledgments

### v2

@@ 4.2 Cell-based search spaces
We now begin our comparative analysis of single and two-stage approaches by applying them to cell-based spaces, which are central in the single-stage NAS literature. We evaluate TangleNAS against DrNAS and SPOS on these spaces. Here, we use the widely studied NAS-Bench-201 (NB201) (dong2020bench) and DARTS (liu-iclr19a) search spaces. We refer the reader to Appendix D for details about these spaces and Appendix F.4 for the experimental setup. We evaluate each method with 4 different random seeds.
Search Type Optimizer Supernet CIFAR-10 (%) ImageNet (%) Search Time (hrs)
Single-Stage DrNAS WS $2.625\pm 0.075$ 26.290 9.1
TangleNAS WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.691}$ 7.4
Two-Stage SPOS+RS WE $2.965\pm 0.072$ 27.114 18.7
SPOS+ES WE $3.200\pm 0.065$ 27.320 14.8
Table 3: Comparison of test errors of single and two-stage methods on the DARTS search space.
Our contribution on these spaces is two-fold. Firstly, we study the effects of weight-entanglement on cell-based spaces in conjunction with single-stage methods. To this end, we entangle the weights of similar operations with different kernel sizes on both search spaces. For NB201, the weights of the $\times1$ and $\times3$ convolutions are entangled, and in the DARTS search space, the weights of dilated and separable convolutions with kernel sizes $\times3$ and $\times5$ are (separately) entangled. Secondly, we study SPOS on the NB201 and DARTS search spaces. To the best of our knowledge, we are the first to study a two-stage method like SPOS in such cell search spaces.
@@
Search Type Optimizer Supernet CIFAR-10 CIFAR-100 ImageNet16-120 Search Time (hrs)
Single-Stage DrNAS WS $\mathbf{94.360\pm 0.000}$ $72.245\pm 0.732$ $\mathbf{46.370\pm 0.000}$ 20.9
TangleNAS WE $\mathbf{94.360\pm 0.000}$ $\mathbf{73.510\pm 0.000}$ $\mathbf{46.370\pm 0.000}$ 20.4
Two-Stage SPOS+RS WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.146$ 29.5
SPOS+ES WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.015$ 26.7
Table 4: Comparison of test-accuracies of single and two-stage methods on NB201 search space.
@@ 4.3 Macro Search Spaces
Given the promising results of TangleNAS on the toy and cell-based spaces, we now extend our evaluation to the home base of two-stage methods. We study TangleNAS on a vision transformer space (AutoFormer-T and -S) and a convolutional space (MobileNetV3), both of which were previously proposed and examined using two-stage methods by chen-iccv21b and cai-iclr20a, respectively. Additionally, we explore a language model transformer search space centered around GPT-2 (radford-openaiblog19a).
@@ 4.3.1 AutoFormer
Search Type Optimizer CIFAR-10 CIFAR-100
Inherit (%) Fine-tune (%) Retrain (%) Params FLOPS Inherit (%) Fine-tune (%) Retrain (%) Params FLOPS
Single-Stage TangleNAS 93.570 $\mathbf{97.702\pm 0.017}$ $\mathbf{97.872\pm 0.054}$ 8.685M 1.946G 76.590 $\mathbf{82.615\pm 0.064}$ $\mathbf{82.668\pm 0.161}$ 8.649M 1.939G
Two-Stage SPOS+RS 94.290 $97.605\pm 0.038$ $97.767\pm 0.024$ 8.512M 1.910G 78.210 $82.407\pm 0.026$ $82.210\pm 0.142$ 8.476M 1.905G
SPOS+ES 94.100 $97.632\pm 0.047$ $97.643\pm 0.023$ 7.230M 1.659G 77.970 $82.517\pm 0.140$ $82.518\pm 0.114$ 8.245M 1.859G
Table 5: Test Accuracies on the AutoFormer-T space for CIFAR-10 and CIFAR-100 (across 4 random seeds)
We evaluate TangleNAS on the AutoFormer-T and AutoFormer-S spaces introduced by chen-iccv21b, based on vision transformers. The search space consists of the choices of embedding dimensions and number of layers, and for each layer, its MLP expansion ratio and the number of heads it uses to compute attention. More details can be found in Table 20 in the appendix. The embedding dimension remains constant across the network, while the number of heads and the MLP expansion ratio change for each layer. This results in a search space of about $10^{13}$ architectures. We train our supernet using the same training hyperparameters and pipeline as used in AutoFormer, and use its evolutionary search as our baseline.
SuperNet-Type NAS Method ImageNet (%) Datasets Params FLOPS
CIFAR-10 (%) CIFAR-100 (%) Flowers (%) Pets(%) Cars(%)
AutoFormer-T SPOS+ES 75.474 98.019 86.369 98.066 91.558 91.935 5.893M 1.396G
AutoFormer-T TangleNAS 78.842 98.249 88.290 98.066 92.347 92.396 8.981M 2.000G
AutoFormer-S SPOS+ES 81.700 99.100 90.459 97.900 94.853 92.545 22.900M 5.100G
AutoFormer-S TangleNAS 81.964 99.120 90.459 98.326 95.070 92.371 28.806M 6.019G
Table 6: Evaluation on the AutoFormer-T space on downstream tasks. ImageNet-1k validation accuracies are obtained through inheritance, whereas the test accuracies for the other datasets are achieved through fine-tuning the ImageNet-pretrained model.
In Tables 5 and 6, we evaluate TangleNAS against AutoFormer on the AutoFormer-T and AutoFormer-S spaces. Interestingly, we observe that although AutoFormer sometimes outperforms TangleNAS upon inheritance from the supernet, the TangleNAS architectures are always better upon fine-tuning and much better upon retraining. For ImageNet-1k we obtain an improvement of $3.368\%$ and $0.264\%$ on AutoFormer-T and AutoFormer-S spaces, respectively (see Table 6).
@@ 4.3.2 MobileNetV3
Next, we study a convolutional search space based on the MobileNetV3 architecture. The search space is defined in Table 19 in the appendix and contains about $2\times 10^{19}$ architectures. This follows from the search space designed by OFA (cai-iclr20a), which searches for kernel size, number of blocks, and channel-expansion factor.
Search Type Optimizer Top-1 acc (%) Params FLOPS
Single-Stage TangleNAS 77.424 7.580M 528.80M
Two-Stage OFA+RS 77.046 6.870M 369.160M
OFA+ES 77.050 7.210M 420.500M
Largest Arch - 77.336 7.660M 566.170M
Table 7: Evaluation on MobileNetV3.
During the supernet training, we activate all choices in our supernet at all times. Table 7 shows that on this OFA search space, TangleNAS outperforms OFA (based on both unconstrained evolutionary and random search on the pre-trained OFA supernet). Notably, TangleNAS even yields an architecture with higher accuracy than the largest architecture in the space (while OFA yields worse architectures).
4.3.3 Language Modeling (LM) Space
Finally, given the growing interest in efficient large language models and recent developments in scaling laws (kaplan2020scaling; hoffmann2022training), we study our efficient and scalable TangleNAS method on a language-model space for two different scales. We create our language model space around a smaller version of nanoGPT model [^1] https://github.com/karpathy/nanoGPT and the model at its original size. In this transformer search space, we search (with 4 random seeds) for the embedding dimension, number of heads, number of layers, and MLP ratios (as defined in Table 18 of the appendix). We weight-entangle all of these by combi-superposition in four dimensions.
Architecture Search-Type Loss Perplexity $\mathbf{\downarrow}$ Params Inference time (s)
GPT-2 Manual 3.077 21.690 123.590M 113.30
TangleNAS Automated 2.904 18.243 116.519M 102.50
(a) Comparison of fine-tuning architecture discovered by TangleNAS and GPT-2 on Shakespeare dataset.
Search Type Optimizer Test loss Perplexity $\mathbf{\downarrow}$ Params Inference Time (s)
Single-Stage TangleNAS $\mathbf{1.412\pm 0.011}$ 4.104 87.010M 93.88
Two-Stage SPOS+RS $1.433\pm 0.005$ 4.191 77.570M 85.17
SPOS+ES $1.444\pm 0.013$ 4.238 78.750M 87.10
(b) Comparison of single and two-stage methods on language model search space. We report the test loss and perplexity on the TinyStories dataset.
Table 8: Evaluation on Language Model Space
For each of these transformer search dimensions, we consider 3 different choices.We train our language models on the TinyStories (eldan-arxiv2023) dataset (for the smaller version of nanoGPT) and on OpenWebText (Gokaslan2019OpenWeb) (for nanoGPT at its original size). Furthermore, we fine-tune the model trained on OpenWebText on the Shakespeare dataset. On the smaller scale, we beat the SPOS+ES and SPOS+RS baselines as presented in Table 8(b). This improvement is statistically very significant with a two-tailed p-value of $0.0064$ for ours v/s SPOS+ES and p-value of $0.0127$ for ours v/s SPOS+RS. As shown in Table 8(a), we obtain a smaller model on OpenWebText while achieving better perplexity after fine-tuning on Shakespeare [^2] https://huggingface.co/datasets/karpathy/tiny_shakespeare. This model is also very efficient during inference time in comparison to GPT-2.
@@ 5 Results and Discussion
For a more thorough evaluation, we now compare different properties of single and two-stage methods, focusing on their any-time performance, the impact of the train-validation split ratios, and the Centered Kernel Alignment (kornblith2019similarity) (see Appendix H) of the supernet feature maps. Additionally, we study the effect of pretraining, fine-tuning and retraining on the AutoFormer space for CIFAR-10 and CIFAR-100, as well as the downstream performance of the best model pre-trained on ImageNet across various classification datasets. We conclude by discussing the insights derived from TangleNAS in designing architectures on real world tasks.
@@
In practice, we observe that vanilla gradient-based NAS methods are memory and compute expensive in comparison to both two-stage methods and our TangleNAS approach with weight-superposition. While the time and space complexity of single-stage methods is $\mathcal{O}(n)$, where $n$ is the number of operation choices, TangleNAS, similar to two-stage methods, maintains $\mathcal{O}(1)$. On the NB201 and DARTS search spaces we observe a 25.28% and 35.54% reduction in memory requirements for TangleNAS over DrNAS with WS. This issue only exacerbates for weight-entangled spaces like AutoFormer, MobileNetV3 and GPT, making the application of vanilla gradient-based methods practically infeasible.
Anytime performance
NAS practitioners often emphasize rapid discovery of competitive architectures. This is especially important given the rising costs of training large and complex neural networks, like Transformers. Thus, strong anytime performance is crucial for the practical deployment of NAS in resource-intensive environments. Therefore, we examine the anytime performance of TangleNAS. Figure 4 demonstrates that TangleNAS (DrNAS+WE) is faster than its baseline method (DrNAS+WS). Similarly, Figure 5 shows that TangleNAS has better anytime performance than AutoFormer.
(a) CIFAR-100
@@
Figure 4: Test accuracy evolution over epochs for NB201.
(a) CIFAR-10
(b) CIFAR-100
Figure 5: Any Time performance curves of AutoFormer vs. Ours.
Effect of fraction of training data
One-shot NAS commonly employs a 50%-50% train-valid split for cell-based spaces and 80%-20% for weight-entangled spaces. To eliminate possible biases, we tested our method on various data splits within each search space. The findings, detailed in Section B, show consistent results across these splits. Specifically, single-stage methods prove to be robust and performant across different training fractions and search spaces, compared to two-stage methods.
@@
5.1 Insights from NAS
@@
In transformer spaces, reducing the MLP-ratio in the initial layers has a relatively low impact on performance (Figure 6) and can often work competitively or outperform handcrafted architectures. This observation is consistent across ViT and Language Model spaces. Conversely, number of heads and embedding dimension have a significant impact.
Pruning a few of the final layers also has a relatively low impact on performance. In the MobileNetV3 space, we find a strong preference for a larger number of channels and larger network depth. In contrast, we discover that scaling laws for transformers may not necessarily apply in convolutional spaces, especially for kernel sizes - the earlier layers favor $\times5$ kernel sizes while later ones prefer $\times7$ (3 being the smallest and 7 the largest).
@@
We examined the effects of inheriting, fine-tuning, and retraining in the AutoFormer space on the CIFAR-10 and CIFAR-100 datasets. We observe that retraining generally surpasses both fine-tuning and inheriting. This raises questions about the correlation between inherited and retrained accuracies of architectures in two-stage methods and the potential training interference identified by xu2022analyzing. A strong correlation is crucial for two-stage methods, which use the performance of architectures with inherited weights as a proxy for true performance in the black-box search. Indeed, while the SPOS+RS and SPOS+ES methods perform well with inherited weights, TangleNAS exceeds their performance after fine-tuning and retraining.
@@
Lastly, we study the impact on fine-tuning the best model obtained from the search on downstream datasets. We follow the fine-tuning pipeline proposed in AutoFormer and fine-tune on different fine- and coarse-grained datasets. We observe from Table 6 that the architecture discovered by TangleNAS on ImageNet is much more performant in fine-tuning to various datasets (CIFAR-10, CIFAR-100, Flowers, Pets and Cars) than the architecture discovered by SPOS.
6 Conclusion and Broader Impact
In this paper, we compare single-stage and two-stage NAS methods, traditionally used for different search spaces, and introduce single-stage NAS to weight-entangled spaces, usually the domain of two-stage methods. We empirically evaluate our single-stage method, TangleNAS, on a diverse set of weight-entangled search spaces and tasks, showcasing its ability to outperform conventional two-stage NAS methods while enhancing search efficiency. Our positive results on macro-level search spaces suggest this approach could advance the development of modern architectures like Transformers within the NAS community. A recent work (klein2023structural) starts training of the supernet from the largest pretrained model, subsequently fine-tuning it. Our method, which now renders single-stage methods applicable to broader families of search spaces (e.g., transformers), can similarly benefit from initialization with pretrained models, achieving additional computational savings. We leave this for future work.
This study addresses the high energy consumption associated with neural architecture search, particularly in black-box techniques that demand extensive computational resources to train many architectures from scratch. Our proposed method falls in the family of one-shot NAS methods, and hence significantly reduces energy usage and identifies efficient architectures with far less energy consumption than manual tuning.
Acknowledgments